• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 12:01:18 on November28,2000

    Entry number 52558

    This entry is a followup to: 52527

    keyword=trigger modifications (cont)

    The trigger modifications on the Right Spectrometer are complete.
    I have checked things out, looks ok. Some notes to add to halog 52527:
    1. The "official" detmap in ~adaq/detmap/detmap.config (which is
    inserted into the datastream and used by various online codes) was
    updated to reflect the fact that the TDC signals for S2 plane on R-arm
    were moved.
    2. New triggers on R-arm, more details
    in ~adaq/e99007/twoarm/trigger/README_e99007
    T1 = (S1-OR).AND.(S2-OR) ...... a simple overlap of S1,S2 planes
    T2 = exclusive of T1, and require 2 out of 3 from S1,S2, Cerenkov
    T7 = The old S-Ray trigger from MLU
    I request to keep about 1 Hz of each of these for studies.
    3. Related to 1 and 2, the xscaler's scaler.config file was updated.